OpenStudy (anonymous):

If the lengths of two sides f a triangle are 2 and 5 , the length of the third side could be 1) 1 2) 6 3) 2 4) 7

OpenStudy (wolf1728):

choice 2 - six

OpenStudy (wolf1728):

The third side must be less than 7 and greater than 3. http://1728.org/trianinq.htm

OpenStudy (anonymous):

Call the 3rd side is c, the other sides are a and b. We have \[c^2 = a^2+ b^2 -2ab \cos \theta =2^2+5^2-2.2.5.\cos \theta =29-20\cos \theta\] \[-1 < \cos \theta < 1 \rightarrow 9 < c^2 < 49 \rightarrow 3 < c < 7\] For your answer, there are 1 option of third side value: 6 Note: \[cos \theta \neq 1\] because it will be not a triangle

OpenStudy (kc_kennylau):

@linh412986 just triangle inequality is enough :)

OpenStudy (kc_kennylau):

@linh412986 sorry for wasting your effort :/

OpenStudy (anonymous):

@kc_kennylau No matter, I think if having time, it is the best way to practice my basic skill of algebra